OpenStudy (anonymous):

You earn $250 per month from your part-time job. You are in a kayaking club that costs $20 per month, and you save at least $100 each month. Write and solve an inequality to find the possible amounts you have left to spend each month.

OpenStudy (anonymous):

Welcome to Openstudy!

OpenStudy (anonymous):

Hi, can u help me with this problem?

OpenStudy (anonymous):

At most 130 dollars: \[s \le 250 - (29 + 100)\] \[s \le 250 - 120 = s \le 130\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Any time! And again welcome to OS! If you need anything please pm me :)

OpenStudy (anonymous):

Ok